OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 3 < -9 or x + 5 greater than or equal to 10 A. A number line with an open circle on -6, shading to the left, and a closed circle on 5, shading to the right. B. A number line with a closed circle on -6, shading to the left, and an open circle on 5, shading to the right. C. A number line with an open circle on -6, a closed circle on 5, and shading in between. D. A number line with a closed circle on -6, an open circle on 5, and shading in between.

OpenStudy (texaschic101):

x - 3 < -9 x < -9 + 3 x < -6 on the second part, is it supposed to be 5x >= 10 ?

OpenStudy (anonymous):

Yes

OpenStudy (texaschic101):

x + 5 >= 10 x >= 10 - 5 x >= 5 ok....we have x is less then -6 and greater then or equal to 5 correct graph is : A

OpenStudy (anonymous):

Thank you so much :) Can i ask you another one please?

OpenStudy (texaschic101):

ask away :)

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 18 less than or greater to 10 and x + 18 greater than or equal to 5 A. A number line with a closed circle on -13, shading to the left, and a closed circle on -8, shading to the right. B. 10 greater than or equal to x + 18 greater than or equal to 5 C. -13 less than or greater to x less than or greater to -8 D. A number line with a closed circle on -13, a closed circle on -8, and shading in between. and this Choose the missing step in the given solution to the inequality -2x - 14 > -4 + 3x. -2x - 14 > -4 + 3x -5x - 14 > -4 -missing step- x < -2 A. -5x < 10 B. -5x > -28 C, x < 2 D. -5x > 10 Thank you so much for the help :)

OpenStudy (texaschic101):

x + 18 <= 10 x + 18 >= 5 x <= 10 - 18 x >= 5 - 18 x <= -8 x >= - 13 so we have x less then or equal to -8 and greater then or equal to -13 -13 <= x <= -8 Not equivalent to A -2x - 14 > -4 + 3x -5x - 14 > -4 -5x > 10 --> missing step x < -2

OpenStudy (anonymous):

Thank you im sure i will have more questions later i will post as a new one

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

wait answer A isnt equvlent or answer C?